# KBQA-o1: Agentic Knowledge Base Question Answering with Monte Carlo Tree Search

### v1

@@ 2 Related Work
Knowledge Base Question Answering. Before the rise of LLMs, the KBQA methods could be divided into information-retrieval-based (IR-based) (Sun et al., 2018, 2019; Zhang et al., 2022) and semantic-parsing-based (SP-based) (Ye et al., 2022; Shu et al., 2022). In the era of LLMs, LLM-based KBQA methods can be divided into two categories: end-to-end and step-by-step. End-to-end methods take advantage of in-context learning (ICL) (Li et al., 2023; Nie et al., 2024) or fine-tuned open-source LLMs (Luo et al., 2024b, a) to equip LLMs with the ability to generate queries. The step-by-step methods (Gu et al., 2023; Sun et al., 2024; Xiong et al., 2024; Huang et al., 2024) follow a reasoning process on the graph to gradually find the answers. In this paper, we propose the first heuristic KBQA method.
LLMs and LLM-powered Agent. LLMs have shown significant advantages in semantic understanding, generation, and reasoning. Powered by CoT(Wei et al., 2022; Li et al., 2024), guiding LLMs to think step by step can further enhance reasoning capabilities. ReAct (Yao et al., 2023b) introduces a prompt-based agent, using tools to interact with the environment. Lightman et al. (2024) verifies the CoT by rewarding the process and the outcome. On the other hand, to expand the thought space in every step, ToT (Yao et al., 2023a) is designed to reason in a tree-like manner. RAP (Hao et al., 2023) employs MCTS, a heuristic algorithm applied in AlphaGo, AlphaGeometry (Trinh et al., 2024) and the OpenAI’s large reasoning model GPT-o1.
@@ 6 Conclusion
This work introduces KBQA-o1, a novel agentic KBQA method with MCTS. However, it has limitations, including the need for more fine-grained policy and reward mechanisms, and challenges in large-scale transfer. To address these, we propose four future directions: (1) exploring reinforcement learning methods such as DPO for continual learning, (2) enriching logical operators for broader reasoning capabilities, (3) adapting the framework to specialized domains like medicine and law, and (4) expanding to multimodal, multilingual, and multi-agent settings, detailed in Appendix K. This work does not raise ethical concerns, as it uses publicly available datasets and focuses on advancing KBQA technology, with positive societal impacts such as supporting decision-making in critical domains.

### v3

@@ 2 Related Work
Knowledge Base Question Answering. Before the rise of LLMs, KBQA methods could be divided into information-retrieval-based (IR-based) (Sun et al., 2018, 2019; Zhang et al., 2022) and semantic-parsing-based (SP-based) (Ye et al., 2022; Shu et al., 2022). In the era of LLMs, LLM-based KBQA methods can be divided into two categories: end-to-end and step-by-step. End-to-end methods take advantage of in-context learning (ICL) (Li et al., 2023; Nie et al., 2024) or fine-tuning (Luo et al., 2024b, a) to enable LLMs to generate queries. The step-by-step methods (Gu et al., 2023; Sun et al., 2024; Huang et al., 2024) follow a reasoning process on the graph to gradually find the answers. In this paper, we propose the first heuristic KBQA method.
LLMs and LLM-powered Agent. LLMs have shown significant advantages in generation and reasoning (Wang et al., 2024). Powered by CoT(Wei et al., 2022; Li et al., 2024), guiding LLMs to think step by step can further enhance reasoning capabilities. ReAct (Yao et al., 2023b) introduces a prompt-based agent, using tools to interact with the environment. Lightman et al. (2024) verifies CoT by rewarding the process and outcome. On the other hand, to expand the thought space in every step, ToT (Yao et al., 2023a) reasons in a tree-like manner. RAP (Hao et al., 2023) employs MCTS, a heuristic algorithm applied in AlphaGo.
@@ 6 Conclusion
Acknowledgments
This work is supported by the National Science Foundation of China (Grant No. 62176026, Grant No. 62473271, and Grant No. 62406036). This work is also supported by the BUPT Excellent Ph.D. Students Foundation (No. CX2023133) and the BUPT Innovation and Entrepreneurship Support Program (No. 2024-YC-A091).
@@ Acknowledgments
This work introduces KBQA-o1, a novel agentic approach to knowledge base question answering (KBQA) using Monte Carlo Tree Search (MCTS). While effective, it still faces limitations in fine-grained policy design and scalability to larger domains. To address these challenges, we outline four future directions: (1) exploring reinforcement learning for continual learning, (2) enhancing logical reasoning capabilities, (3) adapting to specialized domains such as medicine and law, and (4) extending to multimodal, multilingual, and multi-agent settings, as detailed in Appendix K. This work poses no ethical concerns, as it relies solely on publicly available datasets and aims to advance KBQA technology with positive societal impacts.